# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ 3 Theoretical Contribution
Problem Setup. Assume that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. samples drawn from the Laplace gating Gaussian MoE of order $k_{*}$ whose conditional density function $p_{G_{*}}(Y|X)$ is given by:
$p_{G_{*}}(Y|X)=\sum_{i=1}^{k_{*}}$ $\mathrm{softmax}(\mathrm{TopK}(-\|W_{i}^{*}-X\|,K;\beta_{i}^{*}))\cdot f(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*}),$ (4)
where we define for any vectors $v=(v_{i})_{i=1}^{k_{*}}$ and $u=(u_{i})_{i=1}^{k_{*}}$ that $\mathrm{softmax}(v_{i}):=\frac{\exp(v_{i})}{\sum_{j=1}^{k_{*}}\exp(v_{j})}$ and
$\mathrm{TopK}(v_{i},K;u_{i}):=\begin{cases}v_{i}+u_{i},\hskip 23.33147pt\text{if }v_{i}\text{ is in the top }K\text{ elements of }v;\\ -\infty,\hskip 34.14322pt\text{otherwise}.\end{cases}$
Above, $f(\cdot|\mu,\nu)$ denotes a univariate Gaussian density function with mean $\mu$ and variance $\nu$. For ease of the presentation, we denote $G_{*}:=\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{*})\delta_{(W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})}$ as a true but unknown mixing measure associated with unknown parameters $(\beta_{i}^{*},W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})$ for $i\in\{1,2,\ldots,k_{*}\}$. In the paper, we specifically consider two settings of the true number of experts $k_{*}$: (i) Exact-specified setting: when $k_{*}$ is known; (ii) Over-specified setting: when $k_{*}$ is unknown, and we over-specify the model in equation 4 by a Laplace gating MoE model with $k>k_{*}$ experts. However, due to the space limit, we present only the latter setting, and defer the former setting to Appendix J.
Maximum Likelihood Estimation. We use the maximum likelihood method to estimate the unknown mixing measure $G_{*}$. Notably, under the over-specified setting, as the true density $p_{G_{*}}(Y|X)$ is associated with the top $K$ experts which are possibly approximated by more than $K$ fitted experts, we need to select $\overline{K}>K$ experts in the formulation of density estimation to guarantee its convergence to $p_{G_{*}}(Y|X)$. In particular, the MLE is given by
$\widehat{G}_{n}\in\operatorname*{arg\,max}_{G\in\mathcal{G}_{k}(\Theta)}\frac{1}{n}\sum_{i=1}^{n}\log(\overline{p}_{G}(Y_{i}|X_{i})),$ (5)
where $\mathcal{G}_{k}(\Theta):=\{G=\sum_{i=1}^{k^{\prime}}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}:1\leq k^{\prime}\leq k,\ (W_{i},a_{i},b_{i},\nu_{i})\in\Theta\}$ denotes the set of all mixing measures with at most $k$ components and
$\overline{p}_{G}(Y|X)=\sum_{i=1}^{k_{*}}\mathrm{softmax}($ $\mathrm{TopK}(-\|W_{i}-X\|,\overline{K};\beta_{i}))\cdot f(Y|(a_{i})^{\top}X+b_{i},\nu_{i}).$ (6)
Given the MLE defined in equation 5, we first demonstrate in Theorem 3.1 that the convergence rate of density estimation under the Laplace gating Gaussian MoE is parametric on the sample size $n$.
Theorem 3.1 (Density estimation).
The density estimation $\overline{p}_{\widehat{G}_{n}}(Y|X)$ converges to the true density $p_{G_{*}}(Y|X)$ under the Total Variation distance at the following rate:
$\mathbb{E}_{X}[V(\overline{p}_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]=\mathcal{O}(\sqrt{\log(n)/n}).$
Proof of Theorem 3.1 is in Appendix K.3. The parametric rate $\mathcal{O}(\sqrt{\log(n)/n})$ of the conditional density function $\overline{p}_{\widehat{G}_{n}}$ indicates that if there exists a loss function among parameters $\mathcal{D}$ such that $\mathbb{E}_{X}[V(\overline{p}_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}(\widehat{G}_{n},G_{*})$, then we will achieve the parameter and expert estimation rates via the bound $\mathcal{D}(\widehat{G}_{n},G_{*})=\mathcal{O}(\sqrt{\log(n)/n})$.
Voronoi Loss. Following the above implication, we now define a loss function among parameters based on a notion of Voronoi cells as in manole22refined . Given some mixing measure $G$, we distribute its components $\theta_{i}:=(W_{i},a_{i},b_{i},\nu_{i})$ to the following Voronoi cells, which are generated by the components $\theta^{*}_{j}:=(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ of the true mixing measure $G_{*}$:
for any $1\leq j\leq k_{*}$. Note that, the cardinality of the Voronoi cell $\mathcal{A}_{j}$ is exactly the number of fitted components approximating $\theta^{*}_{j}$. For ease of the presentation, let us denote $\Phi_{i\tau_{j}}(\rho_{1},\rho_{2},\rho_{3},\rho_{4}):=\|W_{i}-W_{\tau_{j}}^{*}\|^{\rho_{1}}+\|a_{i}-a_{\tau_{j}}^{*}\|^{\rho_{2}}+|b_{i}-b_{\tau_{j}}^{*}|^{\rho_{3}}+|\nu_{i}-\nu_{\tau_{j}}^{*}|^{\rho_{4}}$, for any $(\rho_{1},\rho_{2},\rho_{3},\rho_{4})\in\mathbb{R}^{4}$. Then, the Voronoi loss function $\mathcal{D}_{2}(G,G_{*})$ used for our analysis under the over-specified setting is given by:
$\max_{\{\tau_{1},\tau_{2},\ldots,\tau_{K}\}\subseteq[k_{*}]}\Bigg{[}\sum_{j=1}^{K}\Big{|}\sum_{i\in\mathcal{A}_{\tau_{j}}}\exp(\beta_{i})-\exp(\beta^{*}_{\tau_{j}})\Big{|}+\sum_{j\in[K]:|\mathcal{A}_{\tau_{j}}|=1}\sum_{i\in\mathcal{A}_{\tau_{j}}}\exp(\beta_{i})\Phi_{i\tau_{j}}(1,1,1,1)$
$\hskip 99.58464pt+\sum_{j\in[K]:|\mathcal{A}_{\tau_{j}}|>1}\sum_{i\in\mathcal{A}_{\tau_{j}}}\exp(\beta_{i})\Phi_{i\tau_{j}}\Big{(}2,2,\bar{r}(|\mathcal{A}_{\tau_{j}}|),\frac{\bar{r}(|\mathcal{A}_{\tau_{j}}|)}{2}\Big{)}\Bigg{]},$ (8)
where the maximum operator included in the above Voronoi loss is to cope with the complex structure of the $\mathrm{TopK}$ function. Additionally, the notation $\bar{r}(|\mathcal{A}_{\tau_{j}}|)$ stands for the minimum value of $r\in\mathbb{N}$ such that the following system of polynomial equations:
$\sum_{i=1}^{|\mathcal{A}_{\tau_{j}}|}\sum_{\begin{subarray}{c}m_{1}+2m_{2}=s,\\ 1\leq m_{1}+m_{2}\leq r\end{subarray}}\frac{q_{3i}^{2}q_{1i}^{m_{1}}q_{2i}^{m_{2}}}{m_{1}!m_{2}!}=0,~{}\text{for each }s=1,2,\ldots,r,$ (9)
does not have any non-trivial solutions for the unknown variables $\{(q_{1i},q_{2i},q_{3i})\}_{i=1}^{|\mathcal{A}_{\tau_{j}}|}$. A solution to the above system is regarded as non-trivial if at least among variables $q_{1i}$ is different from zero, whereas all the variables $q_{3i}$ are non-zero. It is worth noting that the function $\bar{r}(\cdot)$ was previously studied in Ho-Nguyen-Ann-16 to characterize the convergence behavior of parameter estimation under the location-scale Gaussian mixture models. Ho-Nguyen-Ann-16 also gave some specific values of that function, namely $\bar{r}(2)=4$ and $\bar{r}(3)=6$. Meanwhile, they claimed that it was non-trivial to determine the value of $\bar{r}(m)$ when $m\geq 4$, and further techniques should be developed for that purpose. Since Gaussian MoE models are generalization of the Gaussian mixture models, we also involve the function $\bar{r}(\cdot)$ in our convergence analysis. Now, we provide in the following theorem the convergence rate of parameter estimation under the over-specified setting of the Laplace gating Gaussian MoE model (see also Figure 3 for the empirical convergence rates justifying the theoretical rates in Theorem 3.2).
@@
$\mathbb{E}_{X}[V(\overline{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}_{2}(G,G_{*}).$
@@
(i) The parameters $W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i}$ which are fitted by exactly one component, i.e. $|\mathcal{A}^{n}_{i}|:=|\mathcal{A}_{i}(\widehat{G}_{n})|=1$, enjoy the same estimation rate of order $\mathcal{O}(n^{-1/2})$ (up to some logarithmic factor), which match those in nguyen2023demystifying .
@@
Figure 3: Log-log scaled plots illustrating simulation results under the exact-specified (left) and the over-specified settings (right). The orange curves depict the mean discrepancy between the MLE $\widehat{G}_{n}$ and the true mixing measure $G_{*}$, accompanied by error bars signifying two empirical standard deviations. Additionally, the gray dash-dotted line represents the least-squares fitted linear regression line for these data points. See Appendix I for the experimental details.
@@
(ii) The rates for estimating the parameters $W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i}$ which are fitted by more than one component, i.e. $|\mathcal{A}^{n}_{i}|>1$, are no longer homogeneous. On the one hand, the estimation rates for the parameters $b^{*}_{i}$ and $\nu^{*}_{i}$ are of orders $\mathcal{O}(n^{-1/2\bar{r}(|\mathcal{A}^{n}_{i}|)})$ and $\mathcal{O}(n^{-1/\bar{r}(|\mathcal{A}^{n}_{i}|)})$, respectively, both of which are determined by the function $\bar{r}(\cdot)$ and vary with the number of fitted components $|\mathcal{A}^{n}_{i}|$. Those rates are comparable to their counterparts in nguyen2023demystifying . On the other hand, the estimation rates for the gating parameters $W^{*}_{i}$ and the expert parameters $a^{*}_{i}$ are all of order $\mathcal{O}(n^{-1/4})$, which remains constant with respect to the number of fitted components. Meanwhile, those rates in nguyen2023demystifying depend on a different system of polynomial equations from that in equation 9, which are significantly slower.
@@
Advantage of Laplace Gating on FlexiModal Setting  In the standard Softmax gating nguyen2023demystifying , the similarity score is computed as the inner product of a token’s hidden representation and an expert embedding. However, this approach can lead to representation collapse chi2022on , where a subset of experts dominates the decision-making process, resulting in the redundancy of other experts. This issue likely contributes to the slow rates of estimating expert parameters $a^{*}_{i}$ in this setting (see Table 2). By contrast, the Laplace gating function partially alleviates this problem by computing the similarity score as the $L_{2}$-distance between token representations and expert embeddings. This approach is particularly beneficial when dealing with heterogeneous inputs, such as multimodal data or data with missing modalities, as it stabilizes similarity scores and helps prevent load imbalance. The improved estimation rates for expert parameters $a^{*}_{i}$ under the Laplace gating Gaussian MoE, along with our empirical results on multiple large-scale datasets, substantiate these insights.
@@ 4 Experiments
We demonstrate that FuseMoE can provide accurate and efficient predictions when applied to the FlexiModal setting. We tested FuseMoE on a diverse set of benchmarks, including MIMIC-III (johnson2016mimic, ) and MIMIC-IV (johnson2020mimic, ), CMU-MOSI and MOSEI (zadeh2018multi, ), the Physical Activity Monitoring (PAM) dataset (reiss2012introducing, ), and CIFAR-10 (krizhevsky2009learning, ). Compared to CMU-MOSI and MOSEI, the MIMIC ecosystem exhibits irregular and missing modality patterns and includes distinct modalities unlike PAM and CIFAR-10. Evaluating FuseMoE across these diverse datasets provides various empirical insights into critical aspects of our model’s performance. Comprehensive details on the datasets, metrics, parameters, and additional results are thoroughly presented in the Appendices.
@@
Figure 4: (a) The Laplace gating mechanism enhances CIFAR-10 classification when integrated into Vision-MoE (riquelme2021scaling, ). We employed Vision Transformer (ViT) (dosovitskiy2020image, ) and ViT-small as the backbone models and selectively replaced their FFN layers with MoE layers; (b) FuseMoE improves prediction on PAM dataset over baseline time series models; (c) Per-modality routers and the entropy loss $\mathcal{E}$ mitigate the impact of missing modalities.
@@ 4.1 Main Results
We first apply our method to the CMU-MOSI and MOSEI datasets (zadeh2018multi, ), which utilize visual, acoustic, and textual data for sentiment analysis and emotion recognition tasks. Our methodology employs pre-trained T5 (2020t5, ) for text encoding, librosa (mcfee2015librosa, ) for audio feature extraction, and EfficientNet (tan2019efficientnet, ) for video feature encoding. Table 3 details the performance of various router design mechanisms within our MoE architecture, utilizing the Laplace gating function, compared against representative baselines. The baselines include (1) the early fusion method, Tensor Fusion Network (TFN) (zadeh2017tensor, ); (2) the Multimodal Transformer (MulT), which fuses modalities by modeling their interactions (tsai2019multimodal, ); (3) the Multimodal Adaptation Gate (MAG), which focuses on the consistency and differences across modalities (rahman2020integrating, ); and (4) multimodal fusion using standard MoE with the Softmax gating function. Results indicate that employing an MoE backbone—regardless of the gating function chosen or whether utilizing per-modality routers or a joint experts & router configuration—significantly enhances performance on the multimodal task. This improvement is attributed to the MoE’s ability to effectively allocate specific components to handle distinct input modalities, thus better addressing both inter- and intra-modal relationships.
@@
Subsequently, we evaluate our method using the Vision-MoE framework (riquelme2021scaling, ) on the CIFAR-10 classification task (krizhevsky2009learning, ), with results illustrated in Figure 4(a). In this experiment, we selectively replace the FFN layers with an even number in the Vision Transformer (ViT) models with MoE layers. These results, along with Table 3 on the CMU-MOSI and MOSEI datasets comparing Softmax-gating MoE, indicate that the Laplace gating function surpasses the standard Softmax gating function in performance. This outcome is consistent with our theoretical claims.
@@
We then conduct comprehensive evaluations of FuseMoE on MIMIC-IV (johnson2020mimic, ), and the Physical Activity Monitoring (PAM) dataset (reiss2012introducing, ). These datasets feature multiple input modalities, each characterized by varying degrees of irregular sampling or significant levels of missingness. Our tasks of interest for MIMIC datasets include the 48-hour in-hospital mortality prediction (48-IHM), 25-type phenotype classification (25-PHE), and length-of-stay (LOS) prediction. In addition to the previously mentioned baselines, we have incorporated the HAIM method (soenksen2022integrated, ), a data pipeline specifically designed for integrating multimodal data from the MIMIC-IV dataset. We also include the cross-attention combined with irregular sequences modeling approach (MISTS) (zhang2023improving, ). Table 4 shows the outcomes of combining irregular vital signs and clinical notes from the MIMIC-IV dataset. In addition to the commonly used Softmax gating function, we also evaluated the Gaussian gating function (xu1994alternative, ) as a comparative benchmark. The FuseMoE-based methods surpass baselines in most scenarios, often by a non-trivial margin. Furthermore, we observe that HAIM shows considerable efficacy in extracting features from time series, resulting in a strong performance in the 48-IHM and LOS tasks, which are heavily reliant on such data. However, its performance appears more moderate on the 25-PHE task. The PAM dataset captures daily living activities through 17 sensors, with data from each sensor treated as a separate modality. These modalities are individually processed through time-series and irregularity encoders before being integrated into the FuseMoE framework. Our baselines include the Transformer (vaswani2017attention, ), GRU-D (che2018recurrent, ), SeFT (horn2020set, ), a mTAND-only configuration, and IP-Net (shukla2019interpolation, ). We use the Laplace gating and its joint experts & router structure in these experiments. The results in Figure 4(b) have again shown the efficacy of integrating the irregularity encoder with the MoE fusion layer.
4.2 Ablation Studies
Scalability of FuseMoE with Increasing Modalities
Table 4: Comparison of FuseMoE-based methods (gray) and baselines, utilizing vital signs and clinical notes of MIMIC-IV (johnson2020mimic, ). The best results are highlighted in bold font, and the second-best results are underlined. All results are averaged across 5 random experiments.

### v4

@@ 3 Theoretical Contribution
$\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}_{2}(G,G_{*}).$
@@
(i) The parameters $W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i}$ which are fitted by exactly one component, i.e. $|\mathcal{A}^{n}_{i}|:=|\mathcal{A}_{i}(\widehat{G}_{n})|=1$, enjoy the same estimation rate of order $\mathcal{O}(n^{-1/2})$ (up to some logarithmic factor), which match those in Nguyen et al. (2023).
(a) Exact-specified setting
(b) Over-specified setting
@@
Figure 3: Log-log scaled plots illustrating simulation results under the exact-specified (left) and the over-specified settings (right). The orange curves depict the mean discrepancy between the MLE $\widehat{G}_{n}$ and the true mixing measure $G_{*}$, accompanied by error bars signifying two empirical standard deviations. Additionally, the gray dash-dotted line represents the least-squares fitted linear regression line for these data points. Finally, the loss functions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are defined in equations equation 9 and equation 3, respectively. See Appendix I for the experimental details.
@@
(ii) The rates for estimating the parameters $W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i}$ which are fitted by more than one component, i.e. $|\mathcal{A}^{n}_{i}|>1$, are no longer homogeneous. On the one hand, the estimation rates for the parameters $b^{*}_{i}$ and $\nu^{*}_{i}$ are of orders $\mathcal{O}(n^{-1/2\bar{r}(|\mathcal{A}^{n}_{i}|)})$ and $\mathcal{O}(n^{-1/\bar{r}(|\mathcal{A}^{n}_{i}|)})$, respectively, both of which are determined by the function $\bar{r}(\cdot)$ and vary with the number of fitted components $|\mathcal{A}^{n}_{i}|$. Those rates are comparable to their counterparts in Nguyen et al. (2023). On the other hand, the estimation rates for the gating parameters $W^{*}_{i}$ and the expert parameters $a^{*}_{i}$ are all of order $\mathcal{O}(n^{-1/4})$, which remains constant with respect to the number of fitted components. Meanwhile, those rates in Nguyen et al. (2023) depend on a different system of polynomial equations from that in equation 8, which are significantly slower.
@@
Advantage of Laplace Gating on FlexiModal Setting  In the standard Softmax gating Nguyen et al. (2023), the similarity score is computed as the inner product of a token’s hidden representation and an expert embedding. However, this approach can lead to representation collapse Chi et al. (2022); Pham et al. (2024), where a subset of experts dominates the decision-making process, resulting in the redundancy of other experts. This issue likely contributes to the slow rates of estimating expert parameters $a^{*}_{i}$ in this setting (see Table 2). By contrast, the Laplace gating function partially alleviates this problem by computing the similarity score as the $L_{2}$-distance between token representations and expert embeddings. This approach does not inherently favor any expert based on magnitude, unlike inner product which can be biased towards experts with larger norms. The Laplace gating ensures that all experts have a more balanced opportunity to be selected based on how close they are to the token representation. Therefore, Laplace gating is beneficial when dealing with heterogeneous inputs, such as multimodal data, where its feature distributions can be very different across modalities. This is because it can handle these differences without being overly sensitive to the scale and variance of the input features. In addition, it can gracefully degrade in the presence of missing data, rather than causing abrupt changes in gating probabilities that might occur with inner product-based measures. The improved estimation rates for expert parameters $a^{*}_{i}$ under the Laplace gating Gaussian MoE, along with our empirical results on multiple large-scale datasets, substantiate these insights.
@@ 4 Experiments
We demonstrate that FuseMoE can provide accurate and efficient predictions when applied to the FlexiModal setting. We tested FuseMoE on a diverse set of benchmarks, including MIMIC-III (Johnson et al., 2016a) and MIMIC-IV (Johnson et al., 2020), CMU-MOSI and MOSEI (Zadeh et al., 2018), the Physical Activity Monitoring (PAM) dataset (Reiss and Stricker, 2012), and CIFAR-10 (Krizhevsky et al., 2009). Compared to CMU-MOSI and MOSEI, the MIMIC ecosystem exhibits irregular and missing modality patterns and includes distinct modalities unlike PAM and CIFAR-10. Evaluating FuseMoE across these diverse datasets provides various empirical insights into critical aspects of our model’s performance. Comprehensive details on the datasets, metrics, parameters, and additional results are thoroughly presented in the Appendices.
@@
Figure 4: (a) The Laplace gating mechanism enhances CIFAR-10 classification when integrated into Vision-MoE (Riquelme et al., 2021). We employed Vision Transformer (ViT) (Dosovitskiy et al., 2020) and ViT-small as the backbone models and selectively replaced their FFN layers with MoE layers; (b) FuseMoE improves prediction on PAM dataset over baseline time series models; (c) Per-modality routers and the entropy loss $\mathcal{E}$ mitigate the impact of missing modalities.
@@ 4.1 Main Results
We first apply our method to the CMU-MOSI and MOSEI datasets (Zadeh et al., 2018), which utilize visual, acoustic, and textual data for sentiment analysis and emotion recognition tasks. Our methodology employs pre-trained T5 (Raffel et al., 2020) for text encoding, librosa (McFee et al., 2015) for audio feature extraction, and EfficientNet (Tan and Le, 2019) for video feature encoding. Table 3 details the performance of various router design mechanisms within our MoE architecture, utilizing the Laplace gating function, compared against representative baselines. The baselines include (1) the early fusion method, Tensor Fusion Network (TFN) (Zadeh et al., 2017); (2) the Multimodal Transformer (MulT), which fuses modalities by modeling their interactions (Tsai et al., 2019); (3) the Multimodal Adaptation Gate (MAG), which focuses on the consistency and differences across modalities (Rahman et al., 2020); and (4) multimodal fusion using standard MoE with the Softmax gating function. Results indicate that employing an MoE backbone—regardless of the gating function chosen or whether utilizing per-modality routers or a joint experts & router configuration—significantly enhances performance on the multimodal task. This improvement is attributed to the MoE’s ability to effectively allocate specific components to handle distinct input modalities, thus better addressing both inter- and intra-modal relationships.
@@
Subsequently, we evaluate our method using the Vision-MoE framework (Riquelme et al., 2021) on the CIFAR-10 classification task (Krizhevsky et al., 2009), with results illustrated in Figure 4(a). In this experiment, we selectively replace the FFN layers with an even number in the Vision Transformer (ViT) models with MoE layers. These results, along with Table 3 on the CMU-MOSI and MOSEI datasets comparing Softmax-gating MoE, indicate that the Laplace gating function surpasses the standard Softmax gating function in performance. This outcome is consistent with our theoretical claims.
@@
We then conduct comprehensive evaluations of FuseMoE on MIMIC-IV (Johnson et al., 2020), and the Physical Activity Monitoring (PAM) dataset (Reiss and Stricker, 2012). These datasets feature multiple input modalities, each characterized by varying degrees of irregular sampling or significant levels of missingness. Our tasks of interest for MIMIC datasets include the 48-hour in-hospital mortality prediction (48-IHM), 25-type phenotype classification (25-PHE), and length-of-stay (LOS) prediction. In addition to the previously mentioned baselines, we have incorporated the HAIM method (Soenksen et al., 2022), a data pipeline specifically designed for integrating multimodal data from the MIMIC-IV dataset. We also include the cross-attention combined with irregular sequences modeling approach (MISTS) (Zhang et al., 2023). Table 4 shows the outcomes of combining irregular vital signs and clinical notes from the MIMIC-IV dataset. In addition to the commonly used Softmax gating function, we also evaluated the Gaussian gating function (Xu et al., 1994) as a comparative benchmark. The FuseMoE-based methods surpass baselines in most scenarios, often by a non-trivial margin. Furthermore, we observe that HAIM shows considerable efficacy in extracting features from time series, resulting in a strong performance in the 48-IHM and LOS tasks, which are heavily reliant on such data. However, its performance appears more moderate on the 25-PHE task. The PAM dataset captures daily living activities through 17 sensors, with data from each sensor treated as a separate modality. These modalities are individually processed through time-series and irregularity encoders before being integrated into the FuseMoE framework. Our baselines include the Transformer (Vaswani et al., 2017), GRU-D (Che et al., 2018), SeFT (Horn et al., 2020), a mTAND-only configuration, and IP-Net (Shukla and Marlin, 2019). We use the Laplace gating and its joint experts & router structure in these experiments. The results in Figure 4(b) have again shown the efficacy of integrating the irregularity encoder with the MoE fusion layer.
Table 4: Comparison of FuseMoE-based methods (gray) and baselines, utilizing vital signs and clinical notes of MIMIC-IV (Johnson et al., 2020). The best results are highlighted in bold font, and the second-best results are underlined. All results are averaged across 5 random experiments.
@@
4.2 Ablation Studies
Scalability of FuseMoE with Increasing Modalities